The un-damped system has angular frequency w=15Hz.The system is then damped withdamping constant y=5Hz.at what angular frequency must the system be driven in order to achieve the maximum amplitude ?


1
Expert's answer
2021-05-26T15:32:57-0400

"w_d = w \sqrt{1-ξ^2} \\
= \sqrt{w^2-(ξw)^2} \\
= \sqrt{w^2- γ^2} \\
= \sqrt{15^2-5^2} \\
= \sqrt{200} \\
= 14.14 \;Hz"